OpenStudy (anonymous):

The table below represents the distance of a car from its destination as a function of time. Time (hour) x Distance (miles) y 0 900 1 850 2 800 3 750 Part A: What is the y-intercept of the function and what does this tell you about the car? Part B: Calculate the average rate of change of the function represented by the table between x = 1 to x = 3 and what does the average rate represent? Part C: What would be the domain of this function if the car traveled the same rate until it reached its dest

OpenStudy (anonymous):

destination.

OpenStudy (solomonzelman):

I would edit the question. Remove the "10 points" b/c if mod sees your're toasted !

OpenStudy (solomonzelman):

But, anyway, will continue...

OpenStudy (solomonzelman):

You see that after each hour of driving, the car gets 50 miles closer to the destination. Yes ?

OpenStudy (anonymous):

Yes

OpenStudy (solomonzelman):

Lets do Part A. What does the y-intercept tell you. (I am not good at math, but) y-intercept is 0, so the function that you have can also be named "direct variation" As far as the distance to destination goes, that means that you were "right at the start" and no miles ahead of start . (you know what I mean by start, right ?)

OpenStudy (anonymous):

Yeah

OpenStudy (solomonzelman):

For part B, taking the rate of change would be just like finding the slope of a line in this case. y = -50x + 900 ( hence, 1= -50(1)+900=850 2= -50(2)+900=800 3= -50(3)+900=750 and so forth.. ) I don't see how rate of change would be different between x=1 and x=3, or x=0 and x=5, because it's just the slope of the line (that is connections the 2 points of x=1, y= 850 AND x=3 y=750. (I have to give you hints, because I was warned for doing all of asker's work instead of guiding. Sorry)

OpenStudy (solomonzelman):

Alright, the part C. the domain is.... If you were to graph the line, te 1st x-value would be 0 and the last is 9, so the domain is all values btw 0 and 9.

OpenStudy (solomonzelman):

All I can say is that your question is an algebra 1 but in a camouflaged form ;)

OpenStudy (anonymous):

Haha, thanks ! Can you help me with a few more?

OpenStudy (solomonzelman):

Perhaps, but I stink at math :)

OpenStudy (anonymous):

Haha, nah. You don't stink at math. (:

OpenStudy (anonymous):

The table shows the values of y for each value of x. x y 2 1 3 1 4 -3 5 -3 Part A: Does the table represent a relation that is a function? Justify your answer by using the values shown in the table. Part B: The function f(h) shown below represents Kylie's total earning from working in a cafe for h hours of work. f(h) = 15h + 5 Calculate and interpret the meaning of f(8)? Part C: Write an ordered pair to represent the input and output of the function in Part B when Kylie works for r hours.

OpenStudy (anonymous):

x- 2, 3, 4, 5 y- 1, 1, -3, -3

OpenStudy (anonymous):

I just need help with Part C.

OpenStudy (anonymous):

@SolomonZelman